# Hölder curves with exotic tangent spaces

### v2

@@ 6. Hölder parametrizations of carpets $K^{\eta}$
Let $\mathscr{F}_{\infty}=\bigcup_{m\geq 0}\mathscr{F}_{m}$ and let $\mathscr{L}_{\infty}$ be the collection of intervals $I\in\bigcup_{m\geq 0}\mathscr{N}_{m}$ so that $\phi^{\eta}_{\operatorname{{\bf w}}(I)}({\bf 0})$ has an edge emanating upward and an edge emanating to the right in $T_{|\operatorname{{\bf w}}(I)|}$. Note that $\mathscr{L}_{\infty}=\bigcup_{m\geq 0}\{I\in\mathscr{N}_{m}:|\operatorname{{\bf w}}(I)|<m\}$. By (P7) and by design of the families $\mathscr{F}_{m}$, $\mathscr{E}_{m}$, and $\mathscr{N}_{m}$ we have that the restriction of $\zeta$ on $0,1]\setminus\bigcup(\mathscr{F}_{\infty}\cup\mathscr{L}_{\infty})$ is strictly increasing and injective.
@@
To complete the proof, it remains to show that $F$ is $(1/\alpha_{n})$-Hölder continuous. To this end, fix distinct points $x,y\in[0,1]$ and assume that $x<y$, and let $m\geq 0$ be the maximal integer so that there exists an interval $J\in\mathscr{E}_{m}\cup\mathscr{N}_{m}$ with $x,y\in\overline{\zeta(J)}$. The remainder of the proof falls to a case study.
Case 1. If there is an interval $I\in\mathscr{E}_{m+1}(J)$ so that $x$ and $y$ are separated by $\zeta(I)$, then by (P3) and (P4), $|F(x)-F(y)|\leq n^{-m}\sqrt{2}$ and $|x-y|\geq(5n-6)^{-m-2}$, so $|F(x)-F(y)|\leq n^{2}\sqrt{2}|x-y|^{\alpha_{n}}$.
Case 2. If there is an interval $I\in\mathscr{N}_{m+1}(J)\setminus\mathscr{L}_{\infty}$ so that $\zeta(I)$ separates $x$ and $y$, then proceeding similarly as in Case 1. we obtain $|F(x)-F(y)|\leq n^{3}\sqrt{2}|x-y|^{\alpha_{n}}$.
Case 3. If neither Case 1 nor Case 2 holds, then there exist intervals $I_{1},I_{2}\in\mathscr{E}_{m+1}(J)\cup(\mathscr{N}_{m+1}(J)\setminus\mathscr{L}_{\infty})$ so that $x\in\overline{\zeta(I_{1})}$, $y\in\overline{\zeta(I_{2})}$. Then the intersection $\overline{\zeta(I_{1})}\cap\overline{\zeta(I_{2})}=\{p\}$ for some $p$, so we may apply Case 1 to the pairs $x,p$ and $p,y$ to obtain $|F(x)-F(y)|\leq n^{3}\sqrt{2}|x-y|^{\alpha_{n}}$. ∎
@@ 7. Random choice functions
Figure 4. The big triangle represents all words in $\mathcal{A}_{n}^{*}$ that start with $w(\ell-N)$ and have at most $\ell+k+N$ letters. Words of the same the height have the same number of letters. The part of this set where $\eta$ takes the value 2 is colored by red, and the part of the set where $\eta$ takes the value 1 is colored by blue.
@@
Using the notation above, we claim that $\mu_{n}(D^{w})=1$ for $\nu_{n}$-almost every word $w\in\mathcal{A}_{n}^{\mathbb{N}}$. Assuming the claim, by Fubini’s theorem we obtain
@@
For the proof of the claim, recall that
@@
so it suffices to show that the claim holds for every word $w$ in the above set.
@@
so the sets $(E_{m,w})_{m\in\mathbb{N}}$ are independent.
Furthermore, each set $E_{m,w}$ has the same positive measure, so
@@ 8.1. Local cut-points in a class of carpets
We say that $f_{1},\dots,f_{l}$ meet at edges of $0,1]^{2}$ if for every $i\in\{1,\dots,l\}$, there exists $j\in\{1,\dots,l\}\setminus\{i\}$ so that $f_{i}([0,1]^{2})\cap f_{j}([0,1]^{2})$ is an edge of $f_{i}([0,1]^{2})$. Note that necessarily $l\in\{2,3,4\}$.
@@
Recall from the proof of Lemma 4.3 that for any $k\in\mathbb{N}$,
@@
Suppose that $a,b\in E$ are distinct words so that $\phi_{a}^{\eta},\phi_{b}^{\eta}$ do not meet at edges of $0,1]^{2}$. Then $\phi_{a}^{\eta}({\bf 0})-\phi_{b}^{\eta}({\bf 0})\in\{-n^{-m},n^{-m}\}^{2}$. Since $\operatorname{card}(E)\geq 3$, there exists $c\in E\setminus\{a,b\}$ with
@@
Assume now that the claim holds for some $m\in\mathbb{N}$. Let $x\in K^{n,0}$ and $u,v\in\mathcal{A}_{n}^{m+1}$ distinct so that $x\in\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,v}^{2}(K^{n,0})$. By Lemma 8.4, we may assume that $\{w\in\mathcal{A}_{n}^{m+1}:x\in\psi_{n,w}^{2}(K^{n,0})\}=\{u,v\}$. We prove that $\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,w}^{2}(K^{n,0})$ is an edge of $\psi_{n,u}^{2}([0,1]^{2})$.
Let $u=a_{1}\dots a_{m+1}$, $v=b_{1}\dots b_{m+1}$. If $a_{1}=b_{1}$, then applying $(\psi_{n,a_{1}}^{2})^{-1}$ returns us to the claim at $m$, so by the inductive hypothesis we may assume that $a_{1}\neq b_{1}$. Next, if $x$ is not a vertex of $\psi_{n,u}^{2}([0,1]^{2})$, then by Lemma 10.2 the claim holds, so we may further assume that $x$ is a vertex of $\psi_{n,u}^{2}([0,1]^{2})$ and of $\psi_{n,v}^{2}([0,1]^{2})$. If there exists a word $w\in\mathcal{A}_{n}^{m}\setminus\{u(m),v(m)\}$ so that $x\in\psi_{n,w}^{2}(K^{n,0})$, then there exists a $j\in\mathcal{A}_{n}$ with $x\in\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,v}^{2}(K^{n,0})\cap\psi_{n,wj}^{2}(K^{n,0})$, and this is a contradiction.
@@
with $k,k\pm 1\in\{0,1,\dots,n-1\}$. Then there is some $j\in\mathcal{A}_{n}$ so that $\psi_{n,j}^{2}({\bf 0})=(\frac{k}{n},1)$, thus by a straightforward calculation we obtain $x\in\psi_{n,v(m)j}^{2}(K^{n,0})$. As $a_{1}\neq b_{1}$ and necessarily $j\neq b_{m+1}$, $v(m)j\in\mathcal{A}_{n}^{m+1}\setminus\{v,u\}$, and this contradicts our assumption that $\{w\in\mathcal{A}_{n}^{m+1}:x\in\psi_{n,w}^{2}(K^{n,0})\}=\{u,v\}$, concluding the proof. ∎
@@
To prove that $K^{n,0}$ contains no cut-points, fix $y\in K^{n,0}$ and distinct $p,q\in K^{n,0}\setminus\{y\}$. Let $m\in\mathbb{N}$ so that there exist distinct $u,v,w\in\mathcal{A}_{n}^{m}$ for which $\psi_{n,u}^{2}(K^{n,0})$, $\psi_{n,v}^{2}(K^{n,0})$, and $\psi_{n,w}^{2}(K^{n,0})$ are pairwise disjoint with $p\in\psi_{n,v}^{2}(K^{n,0})$, $q\in\psi_{n,u}^{2}(K^{n,0})$, and $y\in\psi_{n,w}^{2}(K^{n,0})$. Then by Lemma 8.5 and by [BV21, Lemma 3.1], there exist distinct $u_{1},\dots,u_{k}\in\mathcal{A}_{n}^{m}$ so that $p\in\psi_{n,u_{1}}^{2}(K^{n,0})$, $q\in\psi_{n,u_{k}}^{2}(K^{n,0})$, and for every $j\in\{1,\dots,k-1\}$, $\psi_{n,u_{j}}^{2}(K^{n,0})\cap\psi_{n,u_{j+1}}^{2}(K^{n,0})$ contains at least two points. Now let $z_{0},\dots,z_{k}\in K^{n,0}$ be distinct points so that $z_{0}=p$, $z_{k}=q$, and $z_{j}\in(\psi_{n,u_{j}}^{2}(K^{n,0})\cap\psi_{n,u_{j+1}}^{2}(K^{n,0}))\setminus\{y\}$ for every $j\in\{1,\dots,k-1\}$ . Then since $\partial[0,1]^{2}\subset K^{n,0}$ by the proof of Lemma 4.3, for every $j\in\{1,\dots,k-2\}$ there is a continuous function $\gamma_{j}:[0,1]\to\psi_{n,u_{j+1}}^{2}(\partial[0,1]^{2})\setminus\{y\}$ so that $\gamma_{j}(0)=z_{j}$, $\gamma_{j}(1)=z_{j+1}$ and necessarily $\gamma_{j}([0,1])\subset K^{n,0}\setminus\{y\}$. By [BV21, Proposition 3.2], there exist continuous $\gamma_{0}:[0,1]\to\psi_{n,u_{1}}^{2}(K^{n,0})$ with $\gamma_{0}(0)=p$, $\gamma_{0}(1)=z_{1}$ and continuous $\gamma_{k-1}:[0,1]\to\psi_{n,u_{k}}^{2}(K^{n,0})$ with $\gamma_{k-1}(0)=z_{k-1}$, $\gamma_{k-1}(1)=q$. Concatenating these curves yields a path inside of $K^{n,0}$ connecting $p$ to $q$ and avoiding $y$, so $y$ cannot be a cut-point for $K^{n,0}$. ∎
@@
By Lemma 4.2, each $\psi_{n,j}^{1}(K^{n,k-1}\cap(0,1)^{2})$ contributes $N_{k-1}$ many local cut-points, and there is one additional local cut-point at $(1/2,2/n)$. Then by the last part of Remark 8.3, every point of $K^{n,k}\setminus(\bigcup_{j\in\mathcal{A}_{n}}\psi_{n,j}^{1}(K^{n,k-1}\cap(0,1)^{2})$ other than $(1/2,2/n)$ is either contained in exactly one image $\psi_{n,v}^{1}(K^{n,0})$ for some $v\in\mathcal{A}_{n}^{k}$ or in an intersection of at least two such images of maps which meet at edges of $0,1]^{2}$, so there are no other local cut-points of $K^{n,k}$ by Corollary 8.7. Therefore, $N_{k}=(5n-6)N_{k-1}+1$. ∎
@@
By (R2), $i_{\ell_{N}-N+1}>(4n-4)$, so $\psi^{\eta(w(\ell_{N}-N))}_{n,i_{\ell_{N}-N+1}}([0,1]^{2})$ does not intersect $\partial[0,1]^{2}$. Thus, for every $N\geq 3$ we get
@@
Now by Remark 2.8, the latter set is equal to $(\phi_{w(\ell_{N})}^{\eta})^{-1}\circ\phi^{\eta}_{w(\ell_{N}+N+k-1)}([0,1]^{2})$, which has diameter no more than $n^{-N-k+2}\sqrt{2}$ and contains $y_{N}$, so $|y_{N}-x_{N}|\leq n^{-N-k+3}$. ∎
In the following lemma, we show that the parts of the sets $X_{N}$ inside of some unit square containing the origin become close to translates of $K^{n,k}$ in terms of Hausdorff distance.
@@
Recall that if $A,B\subset\mathbb{R}^{M}$ are compact sets, then Hausdorff distance is given by
$\operatorname{dist_{H}}(A,B):=\max\left\{\sup_{x\in A}\inf_{y\in B}|x-y|,\sup_{y\in B}\inf_{x\in A}|y-x|\right\},$
and recall from §2.4 that excess is given by $\operatorname{excess}(A,B):=\sup_{x\in A}\inf_{y\in B}|x-y|$. Thus to prove the lemma, it suffices to show
@@
$\operatorname{dist}(K^{\eta}_{w(\ell_{N})},K^{\eta}\setminus K^{\eta}_{w(\ell_{N}-N)})\geq n^{-\ell_{N}}2r.$
@@
Fix $y\in Y_{N}\cap\overline{B}({\bf 0},r)$. If $y\in[0,1]^{2}-y_{N}$, then by Lemma 8.9 there exists some $z\in X_{N}$ so that $|y-z|\leq n^{-N-k+4}$.
@@
By (R2) and since $\phi_{v}^{\eta}(\partial[0,1]^{2})\subset K^{\eta}$ for every finite word $v\in\mathcal{A}_{n}^{*}$ by the proof of Lemma 4.3, we have that
@@
Similarly, we obtain $n^{\ell_{N}}(\phi_{w(\ell_{N})}^{\eta}([0,1]^{2})-x)=[0,1]^{2}-(\phi_{w(\ell_{N})}^{\eta})^{-1}(x)$ and $x_{N}=(\phi_{w(\ell_{N})}^{\eta})^{-1}(x)$. Furthermore, $(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}(x)=(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi_{w(\ell_{N})}^{\eta}(x_{N})$, and so
@@
$y\in(n^{N}(Q-(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi_{w(\ell_{N})}^{\eta}({\bf 0}))-y_{N})\setminus([0,1]^{2}-y_{N}).$
Therefore, there exists $u\in\mathcal{A}_{n}^{2N+k-1}$ so that
@@
Let $r>3$, and recall from the proof of Lemma 8.10 that there exists a minimal $N_{0,r}\in\mathbb{N}$ so that for every integer $N\geq N_{0,r}$,
@@
Furthermore, $N_{0,r}$ is nondecreasing in $r$. By (R1), for each $N\in\mathbb{N}$ there exists a word $v_{N}\in\mathcal{A}_{n}^{N}$ so that $(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi^{\eta}_{w(\ell_{N})}=\psi_{n,v_{N}}^{2}$. Then by Remark 2.8 we obtain
@@
Since $\mathfrak{C}(\mathbb{R}^{2};{\bf 0})$ is sequentially compact in the Attouch-Wets topology, the sequence of sets $(Y_{N}+y_{N})_{N\in\mathbb{N}}$ has a subsequence $(Y_{N_{j}}+y_{N_{j}})_{j\in\mathbb{N}}$ converging to a limit $L$ in the Attouch-Wets topology and so that $(y_{N_{j}})_{j\in\mathbb{N}}$ converges to a point $y_{0}\in[0,1]^{2}$ as $j\to\infty$.
@@
If $j\in\mathbb{N}$ so that $N_{j}\geq N_{0,r+2}$, then
@@
and $D_{r,j}\subset D_{r+2,j}$. Then since for every $N\in\mathbb{N}$, every $v\in\mathcal{A}_{n}^{N}$ has $n^{N}(\psi_{n,v}^{2}({\bf 0})-\psi_{n,v_{N}}^{2}({\bf 0}))\in\mathbb{Z}^{2}$, and since $(Y_{N_{j}}+y_{N_{j}})_{j\in\mathbb{N}}$ converges, there exists $j_{0,r}\in\mathbb{N}$ so that for every integer $j\geq j_{0,r}$, $D_{r,j}=D_{r,j_{0,r}}$. Indeed, if $D_{r,j}\neq D_{r,j+1}$ then
@@
or vice-versa, which cannot happen infinitely often by Lemma 2.4. Consequently, for every $r>3$ and every integer $j\geq j_{0,r+4}$, we have that $(Y_{N_{j}}+y_{N_{j}})\cap\overline{B}({\bf 0},r)=\overline{B}({\bf 0},r)\cap L$, so by Lemma 8.2, Corollary 8.7, and Remark 8.3, $L$ has exactly $\frac{1}{5n-7}((5n-6)^{k}-1)$ many local cut-points inside $0,1]^{2}$ and no other local cut-points. Finally, $L_{k}=L-y_{0}$ is the desired limit.∎

### v3

@@ 6. Hölder parameterizations of carpets $K^{\eta}$
Let $\mathscr{F}_{\infty}=\bigcup_{m\geq 0}\mathscr{F}_{m}$ and let $\mathscr{L}_{\infty}$ be the collection of intervals $I\in\bigcup_{m\geq 0}\mathscr{N}_{m}$ such that $\phi^{\eta}_{\operatorname{{\bf w}}(I)}({\bf 0})$ has an edge emanating upward and an edge emanating to the right in $T_{|\operatorname{{\bf w}}(I)|}$. Note that $\mathscr{L}_{\infty}=\bigcup_{m\geq 0}\{I\in\mathscr{N}_{m}:|\operatorname{{\bf w}}(I)|<m\}$. By (P7) and by design of the families $\mathscr{F}_{m}$, $\mathscr{E}_{m}$, and $\mathscr{N}_{m}$ we have that the restriction of $\zeta$ on $0,1]\setminus\bigcup(\mathscr{F}_{\infty}\cup\mathscr{L}_{\infty})$ is strictly increasing and injective.
@@
To complete the proof, it remains to show that $F$ is $(1/\alpha_{n})$-Hölder continuous. To this end, fix distinct points $x,y\in[0,1]$ and assume that $x<y$, and let $m\geq 0$ be the maximal integer for which there exists an interval $J\in\mathscr{E}_{m}\cup\mathscr{N}_{m}$ with $x,y\in\overline{\zeta(J)}$. The remainder of the proof falls to a case study.
Case 1. If there is an interval $I\in\mathscr{E}_{m+1}(J)$ such that $x$ and $y$ are separated by $\zeta(I)$, then by (P3) and (P4), $|F(x)-F(y)|\leq n^{-m}\sqrt{2}$ and $|x-y|\geq(5n-6)^{-m-2}$, thus $|F(x)-F(y)|\leq n^{2}\sqrt{2}|x-y|^{\alpha_{n}}$.
Case 2. If there is an interval $I\in\mathscr{N}_{m+1}(J)\setminus\mathscr{L}_{\infty}$ such that $\zeta(I)$ separates $x$ and $y$, then proceeding similarly as in Case 1. we obtain $|F(x)-F(y)|\leq n^{3}\sqrt{2}|x-y|^{\alpha_{n}}$.
Case 3. If neither Case 1 nor Case 2 holds, then there exist intervals $I_{1},I_{2}\in\mathscr{E}_{m+1}(J)\cup(\mathscr{N}_{m+1}(J)\setminus\mathscr{L}_{\infty})$ such that $x\in\overline{\zeta(I_{1})}$, $y\in\overline{\zeta(I_{2})}$. Then the intersection $\overline{\zeta(I_{1})}\cap\overline{\zeta(I_{2})}=\{p\}$ for some $p$, thus we may apply Case 1 to the pairs $x,p$ and $p,y$ to obtain $|F(x)-F(y)|\leq n^{3}\sqrt{2}|x-y|^{\alpha_{n}}$. ∎
@@ 7. Random choice functions
Figure 4. The big triangle represents all words in $\mathcal{A}_{n}^{*}$ that start with $w(\ell-N)$ and have at most $\ell+k+N$ letters. Words at the same altitude have the same number of letters. The part of this set where $\eta$ takes the value 2 is colored by red, and the part of the set where $\eta$ takes the value 1 is colored by blue.
@@
Using the notation above, we claim that
(7.1) $\mu_{n}(D^{w})=1,\quad\text{for $\nu_{n}$-almost every word $w\in\mathcal{A}_{n}^{\mathbb{N}}$.}$
Assuming (7.1), by Fubini’s theorem we obtain
@@
For the proof of (7.1), recall that
@@
so it suffices to show (7.1) for every word $w$ in the above set.
@@
hence the sets $(E_{m,w})_{m\in\mathbb{N}}$ are independent.
Furthermore, each set $E_{m,w}$ has the same positive measure, therefore
@@ 8.1. Local cut-points in a class of carpets
We say that $f_{1},\dots,f_{l}$ meet at edges of $0,1]^{2}$ if for every $i\in\{1,\dots,l\}$, there exists $j\in\{1,\dots,l\}\setminus\{i\}$ such that $f_{i}([0,1]^{2})\cap f_{j}([0,1]^{2})$ is an edge of $f_{i}([0,1]^{2})$. Note that necessarily $l\in\{2,3,4\}$.
@@
By Lemma 4.3 and by definition of $K^{n,k}$, we have that for any $k\in\mathbb{N}$,
@@
Suppose that $a,b\in E$ are distinct words such that $\phi_{a}^{\eta},\phi_{b}^{\eta}$ do not meet at edges of $0,1]^{2}$. Then $\phi_{a}^{\eta}({\bf 0})-\phi_{b}^{\eta}({\bf 0})\in\{-n^{-m},n^{-m}\}^{2}$. Since $\operatorname{card}(E)\geq 3$, there exists $c\in E\setminus\{a,b\}$ with
@@
Assume now that the claim holds for some $m\in\mathbb{N}$. Let $x\in K^{n,0}$ and let $u,v\in\mathcal{A}_{n}^{m+1}$ be distinct words such that $x\in\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,v}^{2}(K^{n,0})$. By Lemma 8.4, we may assume that $\{w\in\mathcal{A}_{n}^{m+1}:x\in\psi_{n,w}^{2}(K^{n,0})\}=\{u,v\}$. We prove that $\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,w}^{2}(K^{n,0})$ is an edge of $\psi_{n,u}^{2}([0,1]^{2})$.
Let $u=a_{1}\dots a_{m+1}$, $v=b_{1}\dots b_{m+1}$. If $a_{1}=b_{1}$, then applying $(\psi_{n,a_{1}}^{2})^{-1}$ returns us to the claim at $m$, hence by the inductive hypothesis we may assume that $a_{1}\neq b_{1}$. If $x$ is not a vertex of $\psi_{n,u}^{2}([0,1]^{2})$, then the claim holds by Lemma 2.9. Thus we may assume that $x$ is a vertex of $\psi_{n,u}^{2}([0,1]^{2})$ and of $\psi_{n,v}^{2}([0,1]^{2})$. If there exists a word $w\in\mathcal{A}_{n}^{m}\setminus\{u(m),v(m)\}$ for which $x\in\psi_{n,w}^{2}(K^{n,0})$, then there exists a $j\in\mathcal{A}_{n}$ with $x\in\psi_{n,u}^{2}(K^{n,0})\cap\psi_{n,v}^{2}(K^{n,0})\cap\psi_{n,wj}^{2}(K^{n,0})$, and this is a contradiction.
@@
with $k,k\pm 1\in\{0,1,\dots,n-1\}$. Then there is some $j\in\mathcal{A}_{n}$ such that $\psi_{n,j}^{2}({\bf 0})=(\frac{k}{n},1)$, thus by a straightforward calculation we obtain $x\in\psi_{n,v(m)j}^{2}(K^{n,0})$. As $a_{1}\neq b_{1}$ and necessarily $j\neq b_{m+1}$, $v(m)j\in\mathcal{A}_{n}^{m+1}\setminus\{v,u\}$, and this contradicts our assumption that $\{w\in\mathcal{A}_{n}^{m+1}:x\in\psi_{n,w}^{2}(K^{n,0})\}=\{u,v\}$, concluding the proof. ∎
@@
To prove that $K^{n,0}$ contains no cut-points, fix $y\in K^{n,0}$ and distinct $p,q\in K^{n,0}\setminus\{y\}$. Let $m\in\mathbb{N}$ such that there exist distinct $u,v,w\in\mathcal{A}_{n}^{m}$ for which $\psi_{n,u}^{2}(K^{n,0})$, $\psi_{n,v}^{2}(K^{n,0})$, and $\psi_{n,w}^{2}(K^{n,0})$ are pairwise disjoint with $p\in\psi_{n,v}^{2}(K^{n,0})$, $q\in\psi_{n,u}^{2}(K^{n,0})$, and $y\in\psi_{n,w}^{2}(K^{n,0})$. Then by Lemma 8.5 and by [BV21, Lemma 3.1], there exist distinct $u_{1},\dots,u_{k}\in\mathcal{A}_{n}^{m}$ for which $p\in\psi_{n,u_{1}}^{2}(K^{n,0})$, $q\in\psi_{n,u_{k}}^{2}(K^{n,0})$, and for every $j\in\{1,\dots,k-1\}$, $\psi_{n,u_{j}}^{2}(K^{n,0})\cap\psi_{n,u_{j+1}}^{2}(K^{n,0})$ contains at least two points. Now let $z_{0},\dots,z_{k}\in K^{n,0}$ be distinct points satisfying $z_{0}=p$, $z_{k}=q$, and $z_{j}\in(\psi_{n,u_{j}}^{2}(K^{n,0})\cap\psi_{n,u_{j+1}}^{2}(K^{n,0}))\setminus\{y\}$ for every $j\in\{1,\dots,k-1\}$ . Then since $\partial[0,1]^{2}\subset K^{n,0}$ by Lemma 4.3, for every $j\in\{1,\dots,k-2\}$ there is a continuous function $\gamma_{j}:[0,1]\to\psi_{n,u_{j+1}}^{2}(\partial[0,1]^{2})\setminus\{y\}$ such that $\gamma_{j}(0)=z_{j}$, $\gamma_{j}(1)=z_{j+1}$ and necessarily $\gamma_{j}([0,1])\subset K^{n,0}\setminus\{y\}$. By [BV21, Proposition 3.2], there exist continuous $\gamma_{0}:[0,1]\to\psi_{n,u_{1}}^{2}(K^{n,0})$ with $\gamma_{0}(0)=p$, $\gamma_{0}(1)=z_{1}$ and continuous $\gamma_{k-1}:[0,1]\to\psi_{n,u_{k}}^{2}(K^{n,0})$ with $\gamma_{k-1}(0)=z_{k-1}$, $\gamma_{k-1}(1)=q$. Concatenating these curves yields a path inside of $K^{n,0}$ connecting $p$ to $q$ and avoiding $y$, therefore $y$ cannot be a cut-point for $K^{n,0}$. ∎
@@
By Lemma 4.2, each $\psi_{n,j}^{1}(K^{n,k-1}\cap(0,1)^{2})$ contributes $N_{k-1}$ many local cut-points, and there is one additional local cut-point at $(1/2,2/n)$. Then by the last part of Remark 8.3, every point of $K^{n,k}\setminus(\bigcup_{j\in\mathcal{A}_{n}}\psi_{n,j}^{1}(K^{n,k-1}\cap(0,1)^{2})$ other than $(1/2,2/n)$ is either contained in exactly one image $\psi_{n,v}^{1}(K^{n,0})$ for some $v\in\mathcal{A}_{n}^{k}$ or in an intersection of at least two such images of maps which meet at edges of $0,1]^{2}$, hence there are no other local cut-points of $K^{n,k}$ by Corollary 8.7. Therefore, $N_{k}=(5n-6)N_{k-1}+1$. ∎
@@
By (R2), $i_{\ell_{N}-N+1}>(4n-4)$, hence $\psi^{\eta(w(\ell_{N}-N))}_{n,i_{\ell_{N}-N+1}}([0,1]^{2})$ does not intersect $\partial[0,1]^{2}$. Thus, for every $N\geq 3$ we get
@@
Now by Remark 2.8, the latter set is equal to $(\phi_{w(\ell_{N})}^{\eta})^{-1}\circ\phi^{\eta}_{w(\ell_{N}+N+k-1)}([0,1]^{2})$, which has diameter no more than $n^{-N-k+2}\sqrt{2}$ and contains $y_{N}$, thus $|y_{N}-x_{N}|\leq n^{-N-k+3}$. ∎
In the following lemma, we show that the parts of the sets $X_{N}$ inside of some unit square containing the origin approximate translates of $K^{n,k}$ in terms of Hausdorff distance.
@@
Recall from §2.1 that excess is given by $\operatorname{excess}(A,B):=\sup_{x\in A}\inf_{y\in B}|x-y|$. Thus to prove the lemma, it suffices to show
@@
(8.3) $\operatorname{dist}(K^{\eta}_{w(\ell_{N})},K^{\eta}\setminus K^{\eta}_{w(\ell_{N}-N)})\geq n^{-\ell_{N}}2r.$
@@
Fix $y\in Y_{N}\cap\overline{B}({\bf 0},r)$. If $y\in[0,1]^{2}-y_{N}$, then by Lemma 8.9 there exists some $z\in X_{N}$ such that $|y-z|\leq n^{-N-k+4}$.
@@
By (R2) and since $\phi_{v}^{\eta}(\partial[0,1]^{2})\subset K^{\eta}$ for every finite word $v\in\mathcal{A}_{n}^{*}$ by Lemma 4.3, we have that
@@
Similarly, we obtain $n^{\ell_{N}}(\phi_{w(\ell_{N})}^{\eta}([0,1]^{2})-x)=[0,1]^{2}-(\phi_{w(\ell_{N})}^{\eta})^{-1}(x)$ and $x_{N}=(\phi_{w(\ell_{N})}^{\eta})^{-1}(x)$. Furthermore, $(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}(x)=(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi_{w(\ell_{N})}^{\eta}(x_{N})$, thus
@@
(8.4) $y\in(n^{N}(Q-(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi_{w(\ell_{N})}^{\eta}({\bf 0}))-y_{N})\setminus([0,1]^{2}-y_{N}).$
Therefore, there exists $u\in\mathcal{A}_{n}^{2N+k-1}$ such that
@@
Let $r>3$ and note by definition of $Y_{N}$, by (8.3), and by (8.4) that there exists a minimal $N_{0,r}\in\mathbb{N}$ such that for every integer $N\geq N_{0,r}$,
@@
Furthermore, $N_{0,r}$ is nondecreasing in $r$. By (R1), for each $N\in\mathbb{N}$ there exists a word $v_{N}\in\mathcal{A}_{n}^{N}$ for which $(\phi_{w(\ell_{N}-N)}^{\eta})^{-1}\circ\phi^{\eta}_{w(\ell_{N})}=\psi_{n,v_{N}}^{2}$. Then by Remark 2.8 we obtain
@@
Since $\mathfrak{C}(\mathbb{R}^{2};{\bf 0})$ is sequentially compact in the Attouch-Wets topology, the sequence of sets $(Y_{N}+y_{N})_{N\in\mathbb{N}}$ has a subsequence $(Y_{N_{j}}+y_{N_{j}})_{j\in\mathbb{N}}$ converging to a limit $L$ in the Attouch-Wets topology and further satisfying $(y_{N_{j}})_{j\in\mathbb{N}}$ converges to a point $y_{0}\in[0,1]^{2}$ as $j\to\infty$.
@@
If $j\in\mathbb{N}$ such that $N_{j}\geq N_{0,r+2}$, then
@@
and $D_{r,j}\subset D_{r+2,j}$. Then since for every $N\in\mathbb{N}$, every $v\in\mathcal{A}_{n}^{N}$ has $n^{N}(\psi_{n,v}^{2}({\bf 0})-\psi_{n,v_{N}}^{2}({\bf 0}))\in\mathbb{Z}^{2}$, and since $(Y_{N_{j}}+y_{N_{j}})_{j\in\mathbb{N}}$ converges, there exists $j_{0,r}\in\mathbb{N}$ such that for every integer $j\geq j_{0,r}$, $D_{r,j}=D_{r,j_{0,r}}$. Indeed, if $D_{r,j}\neq D_{r,j+1}$ then
@@
or vice-versa, which cannot happen infinitely often by Lemma 2.3. Consequently, for every $r>3$ and every integer $j\geq j_{0,r+4}$, we have that $(Y_{N_{j}}+y_{N_{j}})\cap\overline{B}({\bf 0},r)=\overline{B}({\bf 0},r)\cap L$, hence by Lemma 8.2, Corollary 8.7, and Remark 8.3, $L$ has exactly $\frac{1}{5n-7}((5n-6)^{k}-1)$ many local cut-points inside $0,1]^{2}$ and no other local cut-points. Finally, $L_{k}=L-y_{0}$ is the desired limit.∎